Please help me solve this
A two digit number is 2 more than 8 times the sum of its digits. Find the number if its ten digit is 6 more than its units digit.

Answer by checkley77(12844)

10X+Y=8(X+Y)+2
X=Y+6
REPLACE X BY (Y+6) IN THE FIRST EQUATION & SOLVE FOR Y
10(Y+6)+Y=8(Y+6+Y)+2
10Y+60+Y=8(2Y+6)+2
11Y+60=16Y+48+2
11Y-16Y=-60+48+2
-5Y=-10
Y=-10/-5
Y=2 ANS. FOR THE DIGITS VALUE.
X=2+6=8 ANS. FOR THE TENS DIGIT.
PROOF:
10*8+2=8(2+8)+2
80+2=6*10+2
82=82